Beale Manufacturing Company has a beta of 1.3, and Foley Industries has a beta of 0.4. The required return on an index fund that holds the entire stock market is 13%. The risk-free rate of interest is 6.5%. By how much does Beale's required return exceed Foley's required return? Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Required return for:

Beale manufacturing company=6.5+1.3*(13-6.5)=14.95%

Foley industries=6.5+0.4*(13-6.5)=9.1%

Hence excess Required return=14.95-9.1

=5.85%